OpenStudy (anonymous):

gary needs to buy fencing to go around his rectangular dog pen, the pen is 8 ft long and has a perimeter of 36 ft. what is the width in ft of the pen? So this is third grade math and I think the answer is 6 but dont know how to explain it to my son.

OpenStudy (anonymous):

the perimeter of a rectangle is 2L+2W with L=length and W=width you know that L=8 so you have 2(8)+2W=36 2x8 is 16 so 16+2W=36 if you minus 16 from each side 16-16+2W=36-16 you get 2W=20 divide each side by 2 2W/2=20/2 W=10ft

OpenStudy (anonymous):

got it?